[Ncep.nhc.nco_contacts] TSB Morning Rounds - Wed Nov 22, 2017

TSB Admin nhc.tsbadmin at noaa.gov
Wed Nov 22 15:48:19 UTC 2017


Here is today's summary of NHC computer operations:
--- Craig Mattocks

----------

TAFB:
1. TSB AWIPS II Sysadmin Chris Mello changed the RWPS download script to  
point to the new location of the ww3 wave model data on the EMC Polar  
server. The data for the Atlantic and Pacific domains are ingesting into  
GFE. TSB will continue to monitor the dataflow.
HSU:

Other items:
1. NCO is still in the process of setting up the chasis for NHC's new blade  
server. The firmware needs to be installed and the network configured so,  
when it arrives at NHC, it can communicate with the VM Center server.

2. Other National Centers reported that the NCEP VPN (secure virtual  
private network) stopped working (temporarily) yesterday afternoon. NCO,  
working with Cisco, discovered an overloaded firewall. They are trying to  
identify the source of the traffic in case the NCEP network is under attack.

3. N-AWIPS dataflows are experiencing delays because the chasis in which  
the NCOSRV3 and NCOSRV4 blade servers reside was misconfigured to route  
traffic through the slow 2G circuit instead of the fast 10G circuit. NCO  
has a fix for the problem, but it will implemented two weeks from now.
NHC RFCs:


-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20171122/2f6fadeb/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list